OpenStudy (anonymous):

Solve the equation using the Zero-Product Property -9n(5n-5)=0 ?

Parth (parthkohli):

The zero product property says that any number multiplied by 0 is zero. Here, as the product is zero, we see that -9n = 0 or 5n - 5 = 0 Let's solve. -9n = 0 n = 0 OR 5n - 5 = 0 5n = 5 n = 1 n = 1 or 0

OpenStudy (anonymous):

thanks! wasn't sure :)

OpenStudy (anonymous):

oops i typed incorrect

Parth (parthkohli):

Haa hehe

OpenStudy (anonymous):

@matricked solve this one? lol (Use the quadratic formula to solve the equation. If necessary, round to the nearest hundredth.) x^2+3=-4 ?

OpenStudy (anonymous):

-4x*

Parth (parthkohli):

Can I solve?

OpenStudy (anonymous):

yes :3

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow x^2 + 4x + 3 = 0 }\) Quadratic formula: \(\Large \color{Black}{\Rightarrow x = {-4 \pm \sqrt{16 + 12} \over 2} }\)

OpenStudy (anonymous):

thats the end result?

Parth (parthkohli):

If you just need the end result, use Wolfram. We're not the ones. Thanks.

OpenStudy (anonymous):

lol well thanks anyway.